### v1

# New entropy, thermodynamics of apparent horizon and cosmology

Abstract
Here, we consider new nonadditive entropy of the apparent horizon $S_{K}=S_{BH}/(1+\gamma S_{BH})$ with $S_{BH}$ being the Bekenstein–Hawking entropy. This is an alternative of the Rényi and Tsallis entropies, that allow us by utilising the holographic principle to develop of a new model of holographic dark energy. When $\gamma\rightarrow 0$ our entropy becomes the Bekenstein–Hawking entropy $S_{BH}$. The generalised Friedmann’s equations for Friedmann–Lemaître–Robertson–Walker (FLRW) space-time for the barotropic matter fluid with $p=w\rho$ were obtained. We compute the dark energy pressure $p_{D}$, density energy $\rho_{D}$ and the deceleration parameter corresponding to our model. From the second modified Friedmann’s equation a dynamical cosmological constant was obtained. We show that at some model parameters $w$ and $\beta$ there are two phases, universe acceleration and deceleration or the phase of the eternal inflation. Thus, our model, by virtue of the holographic principle, can describe the universe inflation and late time of the universe acceleration. The holographic dark energy model with the generalised entropy of the apparent horizon can be of interest for new cosmology.
I Introduction
To describe the current accelerating universe one can introduce in the Einstein equations the cosmological constant. It plays the role of dark energy leading to standard cosmology. Another way to describe acceleration of universe is to explore thermodynamics of apparent horizon Akbar ; Cai ; Cai0 ; Paranjape ; Sheykhi ; Cai1 ; Wang ; Jamil ; Gim ; Fan ; Agostino ; Sanchez because there is a correspondence between gravity and thermodynamics. This is based on the fact that as was shown in Refs. Bekenstein ; Hawking that in black holes, entropy is connected with the horizon area and temperature is linked with the surface gravity Jacobson ; Padmanabhan ; Padmanabhan1 ; Hayward . Also, Friedmann’s equations may be obtained with the help of the first law of apparent horizon thermodynamics. As a result, the apparent horizon of the FLRW space-time is represented as a thermodynamic system and it is a causal boundary Hayward ; Hayward1 ; Bak , and on this boundary the thermodynamics laws are satisfied Akbar ; Cai0 . In flat FLRW universe the apparent horizon is equal to the Hubble horizon. Different entropies Tsallis ; Ren ; Renyi ; Jahromi ; Barrow ; Kaniadakis ; Mejrhit ; Majhi were introduced, due to the long-range nature of gravity, that lead to the generalized Friedmann equations. Motivated by success of this approach, it is of interest to consider other entropies to describe accelerating universe. Because entropies are sources of holographic energy densities they can describe the dark energy of the universe Pavon ; Landim . In this letter, we consider new apparent horizon entropy $S_{K}=S_{BH}/(1+\gamma S_{BH})$ where $S_{BH}$ is the Bekenstein–Hawking entropy and it is a non-extensive entropy measure. When the Bekenstein–Hawking entropy becomes zero our entropy $S_{K}$ also vanishes, $S_{K}$ is the monotonically increasing function of $S_{BH}$ and is positive. As $\gamma\rightarrow 0$ we come to the Bekenstein–Hawking entropy. Here, we consider the case with equation of state (EoS) for barotropic perfect fluid, $p=w\rho$. We will obtain modified Friedmann’s equations with dynamical cosmological constant that leads to dark energy and the universe inflation. Our new holographic model with the non-additive entropy and non-extensive feature of space-time shows the possibility of the universe to accelerate in accordance with observations.
We assume units with $\hbar=c=k_{B}=1$.
II New entropy
We start with new entropy
$S_{K}=-\sum_{i=1}^{W}\frac{p_{i}\ln{p_{i}}}{1-\gamma\ln{p_{i}}}),$ (1)
where $W$ is a number of states. Each state possesses a probability $p_{i}$ with the probability distribution $\{p_{i}\}$ and $\gamma$ is a free parameter. In Eq. (2.1) the summation is performed over all possible system microstates. When $\gamma=0$ entropy (2.1) is converted into the Gibbs entropy
$S_{G}=-\sum_{i=1}^{W}p_{i}\ln(p_{i}).$ (2)
If one assumes that each microstate is populated with equal probability, then $1/p_{i}=W$ ($i=1,2,...,W$) and Eq. (2.2) is converted into the Boltzmann entropy $S_{B}=\ln(W)$. Making use of $1/p_{i}=W$, we obtain from Eq. (2.1)
$S_{K}=\frac{\ln(W)}{1+\gamma\ln(W)}.$ (3)
The Bekenstein–Hawking entropy $S_{BH}=\ln(W)$. Then from Eq. (2.3) one finds
$S_{K}=\frac{S_{BH}}{1+\gamma S_{BH}}.$ (4)
Making use of Eq. (2.4), at $\gamma=0$, we obtain the Bekenstein–Hawking entropy $S_{BH}$. If A and B are two probabilistically independent systems, we have $p^{A+B}_{ij}=p^{A}_{i}p^{B}_{j}$ and one finds nonadditive entropy, $S_{K}(A+B)\neq S_{K}(A)+S_{K}(B)$.
III Apparent horizon thermodynamics
Let us consider the FLRW flat universe which is described by the metric
$ds^{2}=-dt^{2}+a(t)^{2}(dr^{2}+r^{2}d\Omega_{2}^{2}).$ (5)
Here, $a(t)$ is a scale factor and $d\Omega_{2}^{2}$ is the line element of an 2-dimensional unit sphere. The radius of the apparent horizon $R_{h}=a(t)r$, in the FLRW universe, is given by
$R_{h}=\frac{1}{H},$ (6)
where the Hubble parameter of the universe is $H=\dot{a}(t)/a(t)$, with dot over $a(t)$ being the derivative with respect to the cosmological time $t$. The first law of apparent horizon thermodynamics is formulated as
$dE=-T_{h}dS_{h}+WdV_{h},$ (7)
where $W$ is the work density and $E$ is the total energy inside the space which is given by
$E=\rho V_{h}=\frac{4\pi}{3}\rho R_{h}^{3}.$ (8)
The change of the energy inside the apparent horizon is $-dE$. In Eq. (3.4) $\rho$ means the energy density of matter fields and the work density is Hayward ; Hayward1 ; Bak
$W=-\frac{1}{2}\mbox{Tr}(T^{\mu\nu})=\frac{1}{2}(\rho-p),$ (9)
and $p$ is the matter pressure. In our case the horizon entropy is $S_{h}=S_{K}$. The apparent horizon temperature is defined as
$T_{h}=\frac{H}{2\pi}\left|1+\frac{\dot{H}}{2H^{2}}\right|.$ (10)
Making use of first law of apparent horizon thermodynamics (3.3), and by combining Eqs. (3.4) and (3.5) with (3.6) , one finds
$\frac{H}{2\pi}\left|1+\frac{\dot{H}}{2H^{2}}\right|dS_{h}=-\frac{4\pi}{3H^{3}}d\rho+\frac{2\pi(\rho+p)}{H^{4}}dH.$ (11)
With the help of the continuity equation
$\dot{\rho}=-3H(\rho+p),$ (12)
and Eq. (3.7) we obtain
$\frac{H}{2\pi}\left|1+\frac{\dot{H}}{2H^{2}}\right|\dot{S}_{h}=-\frac{4\pi\dot{\rho}}{3H^{3}}\left(1+\frac{\dot{H}}{2H^{2}}\right).$ (13)
IV Generalized Friedmann’s equations
Utilizing Eqs. (3.8), (3.9) and assuming that $01+\dot{H}/(2H^{2})>$ 0 we obtain
$\frac{H}{2\pi}\dot{S}_{h}=\frac{4\pi(\rho+p)}{H^{2}}.$ (14)
By using our entropy function
$S_{h}=S_{K}=\frac{S_{BH}}{1+\gamma S_{BH}},$ (15)
where $S_{BH}=\pi R^{2}_{h}/G=\pi/(GH^{2})$, we obtain from Eq. (4.1) the generalized Friedmann equation
$\frac{\dot{H}}{(1+\gamma\pi/(GH^{2}))^{2}}=-4\pi G(\rho+p).$ (16)
At $\gamma=0$ in Eq. (4.3), we come to the first Friedmann equation for flat universe within Einstein gravity. Making use of Eq. (3.8) and after integration of Eq. (4.3), we find the second generalized Friedmann equation
$H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)=\frac{8\pi G}{3}\rho,$ (17)
where $b=\pi\gamma/G$ and we have used the integration constant $C=2b\ln(b)$. At $\gamma=0$ ($b=0$) one obtains from Eq. (4.4) the Friedmann equation of general relativity. We can represent Eq. (4.4) as follows:
$H^{2}=\frac{8\pi G}{3}\rho+\frac{\Lambda_{eff}}{3}.$ (18)
with
$\Lambda_{eff}=\frac{3b^{2}}{H^{2}+b}+6b\ln\left(\frac{H^{2}+b}{b}\right).$ (19)
The $\Lambda_{eff}$ represents the dynamical (effective) cosmological constant. We depicted the dynamical cosmological constant $\Lambda_{eff}$ versus $H$ at $b=1,2,3$ in Fig. 1.
Figure 1: The function $\Lambda_{eff}$ versus $H$ at $b=\pi\gamma/G=1,2,3$. Figure 1 shows that $\Lambda_{eff}$ increases as $b$ increases. At $H\rightarrow 0$ ($R_{h}\rightarrow\infty$) we have $\Lambda_{eff}\rightarrow 3b$.
According to Fig. 1 when $b$ increases, at fixed $H$, the dynamical cosmological constant $\Lambda_{eff}$ also increases and $\lim_{H\rightarrow 0}\Lambda_{eff}=3b$. By virtue of Eq. (4.5) one finds the density of dark energy
$\rho_{D}=\frac{3b}{8\pi G}\left[\frac{b}{H^{2}+b}+2\ln\left(\frac{H^{2}+b}{b}\right)\right].$ (20)
Defining the normalized density parameters $\Omega_{m}=\rho/(3M_{P}^{2}H^{2})$ and $\Omega_{D}=\rho_{D}/(3M_{P}^{2}H^{2})$, where $M_{P}=1/\sqrt{8\pi G}$ is the reduced Plank mass, one obtains from Eqs. (4.4) and (4.7) that $\Omega_{m}+\Omega_{D}=1$. From Eq. (4.4) one obtains the normalized density for the matter $w=0$
$\Omega_{m}=1-\frac{b^{2}}{(H^{2}H^{2}+b)}-2\frac{b}{H^{2}}\ln\left(\frac{H^{2}+b}{b}\right).$
The plot of $\Omega_{m}$ versus $H$ is depicted in Fig. 2.
Figure 2: The function $\Omega_{m}$ versus $H$ at $b==0.15,0.2,0.3$. Figure 2 shows that $\Omega_{m}$ increases as $b$ increases. As $H\rightarrow\infty$ ($R_{h}\rightarrow 0$) we have $\Omega_{m}\rightarrow 1$.
According to Fig. 2 for the current era $\Omega_{m}=0.26$ that can be realized for some parameter $b$. We will imply here that there is no mutual interaction between various components of the cosmos, from ordinary conservation law, one obtains
$p_{D}=-\frac{\dot{\rho}_{D}}{3H}-\rho_{D}.$ (21)
By virtue of Eqs. (4.7) and (4.8) we find the pressure corresponding to the dark energy
$p_{D}=-\frac{b(b+2H^{2})\dot{H}}{4\pi G(b+H^{2})^{2}}-$
$\frac{3b}{8\pi G}\left[\frac{b}{H^{2}+b}+2\ln\left(\frac{H^{2}+b}{b}\right)\right].$ (22)
Making use of Eqs, (4.3), (4.4) and (4.9) we obtain
$p_{D}=\frac{3b(b+2H^{2})(1+w)}{8\pi GH^{4}}\biggl{[}H^{2}-\frac{b}{H^{2}+b}-$
$2\ln\left(\frac{H^{2}+b}{b}\right)\biggr{]}-\frac{3b}{8\pi G}\left[\frac{b}{H^{2}+b}+2\ln\left(\frac{H^{2}+b}{b}\right)\right].$ (23)
With the help of Eqs. (4.7) and (4.10) one finds EoS for dark energy
$w_{D}=\frac{p_{D}}{\rho_{D}}=$
$\frac{(b+2H^{2})(1+w)}{H^{4}}\left[\frac{H^{2}(H^{2}+b)}{b+2(H^{2}+b)\ln\left(\frac{H^{2}+b}{b}\right)}-1\right]-1.$ (24)
The plot of $w_{D}$ versus $H$ is depicted in Fig. 3.
Figure 3: Left panel: The function $w_{D}$ versus $H$ at $b=1$, $w=0,1/3,2/3$, $G=1$, $\rho_{0}=a_{0}=1$. According to Fig. 3 the $w_{D}$ increases as EoS parameter for the matter $w$ increases, at fixed $H$. At large $H$ EoS parameter for dark energy $w_{D}$ approaches to $-1$. Right panel: In accordance with figure when parameter $b$ increases, at fixed $H$, EoS parameter for dark energy $w_{D}$ also increases and $\lim_{H\rightarrow\infty}w_{D}=-1$.
In accordance with Fig. 3 (Left panel) at $b=1$ and $w=0,1/3,2/3$ when $w$ increases, at fixed $H$, EoS parameter for dark energy $w_{D}$ also increases. Figure 3 shows (Right panel) that at $w=1$ and $b=2,3,4$ if $b$ increases, at fixed $H$, $w_{D}$ aolso increases. It follows from Eq. (4.11) that $\lim_{H\rightarrow\infty}w_{D}=-1$. Thus, dynamical cosmological constant leads to dark energy EoS $w_{D}=-1$ at large Habble parameter $H$.
The second law of thermodynamics requires that $\dot{S}_{K}\geq 0$. Then from Eq. (3.2) one obtains $\dot{S}_{BH}/(1+\gamma S_{BH})^{2}\geq 0$ and this leads to $\dot{S}_{BH}=-2\pi\dot{H}/(GH^{3})\geq 0$. As a result, we have the same requirement as for the Bekenstein–Hawking entropy. For positive Hubble parameter one obtains $\dot{H}\leq 0$. Then from Eq. (3.3) we find $\rho+p\geq 0$ and for $\rho>0$ the EoS parameter $w\geq-1$. It is convenient to use the redshift $z=a_{0}/a(t)-1$, where $a_{0}$ corresponds to a scale factor at the current time. With the aid of continuity equation (3.8) and EoS $p=w\rho$ one obtains the matter density energy
$\rho=\rho_{0}\left(\frac{1+z}{a_{0}}\right)^{3(1+w)},$ (25)
with $\rho_{0}$ being the density energy of matter at the present time. From Eqs. (4.4) and (4.12) we find equation as follows:
$H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)=\frac{8\pi G\rho_{0}}{3}\left(\frac{1+z}{a_{0}}\right)^{3(1+w)}.$ (26)
The Hubble parameter versus redshift $z$ for $G=1$, $\rho_{0}=a_{0}=1$ is plotted in Fig. 4.
Figure 4: Left panel: The function $H$ versus $z$ at $b=0.1$, $w=1/3,0,-0.1$, $G=1$, $\rho_{0}=a_{0}=1$. According to Fig. 4 $H$ increases as $z$ increases. At fixed $H$, when EoS parameter for the matter $w$ increases the redshift $z$ decreases. Right panel: In accordance with figure if parameter $b$ increases at fixed $z$ the Habble parameter also increases.
When redshift $z$ increases the Habble parameter also increases. According to Fig. 4 (Left panel) if parameter $w$ increases the redshift $z$ decreases. Figure 4 (Right panel) shows that when parameter $b$ increases at fixed z the Habble parameter also increases The deceleration parameter is given by
$q=-\frac{\ddot{a}a}{\dot{a}^{2}}=-1-\frac{\dot{H}}{H^{2}}.$ (27)
If $q<0$ we have the acceleration phase of the universe and when $q>0$ the universe decelerates. Making use of Eqs. (4.3), (4.12) and (4.14) we find
$q=\frac{4\pi G\rho_{0}(1+w)(H^{2}+b)^{2}}{H^{6}}\left(\frac{1+z}{a_{0}}\right)^{3(1+w)}-1.$ (28)
Equation (4.15) defines the dependence of the deceleration parameter $q$ on redshift $z$. By virtue of Eqs. (4.13) and (4.15) we obtain the function of deceleration parameter $q$ on $H$
$q=-1+$
$\frac{3(1+w)(H^{2}+b)^{2}}{2H^{6}}\left(H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)\right).$ (29)
We depicted the deceleration parameter $q$ versus the Hubble parameter $H$ for $\rho_{0}=a_{0}=1$ in Fig. 5.
Figure 5: According to Fig. 5 when deceleration parameter $q$ increases the Habble parameter $H$ also increases. Left panel: The deceleration parameter $q$ versus Habble parameter $H$ at $b=1$, $w=0,1/3$, 2/3, $\rho_{0}=a_{0}=1$. When EoS parameter for the matter $w$ increases, at fixed $b$ and $H$, the deceleration parameter $q$ also increases. There are two phases: universe acceleration $q<0$ and deceleration $q>0$. Right panel: in accordance with figure, when parameter $b$ increases, at fixed $w$ and $H$, the deceleration parameter $q$ decreases. Here we have two phases, universe acceleration and deceleration.
Figure 5 shows that there are two phases: universe acceleration and deceleration. Making use of Eq. (4.16) we find the asymptotic
$\lim_{H\rightarrow\infty}q=\frac{3w+1}{2}.$ (30)
Equation (4.17) shows that the asymptotic of the deceleration parameter as $H\rightarrow\infty$ ($R_{h}\rightarrow 0$) does not depend on the entropy parameter $\gamma$. It follows from Eq. (4.17) that when $w>-1/3$ ($q>0$) we have two phases: universe acceleration and deceleration. If $w<-1/3$ the eternal inflation takes place. By virtue of Eq. (4.13) we obtain the redshift $z$ as a function of Habble parameter
$z=-1+$
$a_{0}\left[\frac{3}{8\pi\rho_{0}G}\left(H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)\right)\right]^{1/(3(1+w))}.$ (31)
From Eqs. (4.16) and (4.18) we found the approximate real and positive solutions for the transition redshifts $z_{t}$ (from acceleration to deceleration phases) at $q=0$, $G=1$, $w=0$ (for the matter dominated era) presented in Table 1.
Table 1: The approximate solutions to Eqs. (3.16) and (3.18) for the transition redshifts $q=0$ at $G=1$, $a_{0}=\rho_{0}=1$, $w=0$.
$b$ 1.5 2 2.5 3 3.5 4 4.5 5 5.5 6 6.5
$H$ 4.26 4.92 5.50 6.03 6.51 6.96 7.38 7.78 8.16 8.53 8.88
$z_{t}$ 0.07 0.18 0.27 0.35 0.42 0.49 0.55 0.60 0.65 0.70 0.75
According to Table I when parameter $b$ increases (at fixed $w$) the Hubble parameter $H$ and reshift $z$ also increase for a point $q=0$ between universe acceleration and deceleration. From Eqs. (4.16) and (4.18) for the current era $z=0$, we find solutions for the Habble parameter $H$ and deceleration parameter $q$ for varies $b$, given in Table II.
Table 2: The solutions (approximate) to Eqs. (3.16) and (3.18) for the current era $z=0$ at $G=1$, $a_{0}=\rho_{0}=1$, $w=-2/3$.
$b$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
$H$ 3.05 3.15 3.25 3.33 3.40 3.47 3.54 3.60 3.66
$q$0.54 -0.56 -0.58 -0.59 -0.61 -0.62 -0.63 -0.64 -0.65
In Table II the deceleration parameter $q$ is negative, and therefore, we have here the acceleration phase in the current time. The deceleration parameter at the current era is $q_{0}\approx-0.6$ Roos corresponding to parameter $b\approx 0.5$. By virtue of (3.16) we plotted the Habble parameter $H$ versus redshift $z$ in Fig. 6.
Figure 6: Left panel: The Habble parameter $H$ versus $z$ at $b=0.1$, $w=-0.3,1/3$, 2/3, $G=1$, $\rho_{0}=a_{0}=1$. Figure 6 shows that when $z$ increases, $H$ also increases. When EoS parameter $w$ increases, at fixed $b$, the Habble parameter $H$ decreases. Right panel: When parameter $b$ increases, at fixed $w$, the Habble parameter $H$ also increases.
In accordance with Fig. 6, when redshift $z$ increases, the Habble parameter $H$ also increases. When EoS parameter $w$ increases, at fixed $b$, the Habble parameter $H$ also increases. Figure 6 shows that if parameter $b$ increases, at fixed $w$, the Habble parameter $H$ also increases. Making use of Eqs. (4.16) and (4.18) we depicted the deceleration parameter $q$ versus redshift $z$ in Fig. 7.
Figure 7: Left panel: The deceleration parameter $q$ versus $z$ at $b=0.1$, $w=-0.3,0$, 1/3, $G=1$, $\rho_{0}=a_{0}=1$. In accordance with Figure 7, when $z$ increases the deceleration parameter $q$ decreases. When EoS parameter for the matter $w$ increases, at fixed $b$, the deceleration parameter $q$ also increases. At $w=-0.3$ there is two phases, universe acceleration $q<0$ and deceleration $q>0$ and at $w=0$ and $1/3$ we have only the deceleration phase. Right panel: According to Fig. 5 when parameter $b$ increases, at fixed $w$, the deceleration parameter $q$ also increases. We have here two phases: acceleration and deceleration for $b=0.4$ and $0.6$. For $b=0.2$ one has only the deceleration phase.
Figure 7 shows that when $z$ increases $q$ also increases. When EoS parameter for the matter $w$ increases, at fixed $b$ (Left panel), the deceleration parameter $q$ also increases. At $w=-0.3$ there are two phases, universe acceleration $q<0$ and deceleration $q>0$ but at $w=0$ and $w=1/3$ one has only the deceleration phase. According to Fig. 5 (Right panel), when parameter $b$ increases at fixed $w$ the deceleration parameter $q$ decreases. One has here two phases, acceleration and deceleration for $b=0.4$ and $b=0.6$ but for $b=0.2$ we have the deceleration phase.
V Conclusion
Considering a novel entropy $S_{K}=S_{BH}/(1+\gamma S_{BH})$ we have a property similar to the Bekenstein–Hawking entropy $S_{BH}$. By using our entropy and by employing the holographic principle, we have obtained a dark energy model. The $S_{K}$ vanishes when the apparent horizon radius $R_{h}$ is zero and it increases monotonically when the apparent horizon radius $R_{h}$ increases. The barotropic perfect fluid and flat FLRW universe were studied. We have obtained, from first law of apparent horizon thermodynamics, the modified Friedmann’s equations. The second Friedmann’s equation possesses an addition term corresponding to the density of dark energy which can be treated as a dynamical cosmological constant. Implying that there is no interaction between various components of cosmos, and the dark energy density and pressure $\rho_{D}$ and $p_{D}$ obeys ordinary conservation law, we computed $\rho_{D}$ and $p_{D}$. The EoS parameter for dark energy $w_{D}=p_{D}/\rho_{D}$ was calculated and it was shown that $\lim_{H\rightarrow\infty}w_{D}=-1$. The universe can have two phases, acceleration and deceleration, that is due to holographic dark energy. It should be noted that Barrow and Tsallis entropies also lead to cosmology which ia due to Einsten’s equations with the dynamical cosmological constant Gennaro . We computed the deceleration parameter that shows the possibility, for some model parameters, to describe inflation and acceleration at current era. The transition redshifts (when $q=0$) was calculated (given in Table I) for some parameters. We show (see Table II) that at $b\approx 0.5$ and $w=-2/3$ we have the deceleration parameter $q_{0}\approx-0.6$ which was observed in universe at the current era Roos . Thus, our approach, based on new entropy and leading to modified Friedmann equations, can description universe inflation and late time universe acceleration. It is worth mentioning that the results obtained here are similar to those in another model with different entropy Kruglov .
Data availability statement: No data was used for the research described in the article.

### v3

@@
 Here, we consider new nonadditive entropy of the apparent horizon $S_{K}=S_{BH}/(1+\gamma S_{BH})$ with $S_{BH}$ being the Bekenstein–Hawking entropy.
-This is an alternative of the Rényi and Tsallis entropies, that allow us by utilising the holographic principle to develop of a new model of holographic dark energy.
+This is an alternative of the Rényi and Tsallis entropies, that allow us by utilising the holographic principle to develop of a new model of entropic (holographic) dark energy.
 When $\gamma\rightarrow 0$ our entropy becomes the Bekenstein–Hawking entropy $S_{BH}$.
@@
 From the second modified Friedmann’s equation a dynamical cosmological constant was obtained.
-We show that at some model parameters $w$ and $\beta$ there are two phases, universe acceleration and deceleration or the phase of the eternal inflation.
+We show that at some model parameters $w$ and $\gamma$ there are two phases, universe acceleration and deceleration or the phase of the eternal inflation.
 Thus, our model, by virtue of the holographic principle, can describe the universe inflation and late time of the universe acceleration.
+It is shown that entropic cosmology with our entropy proposed is equivalent to cosmology based on the teleparallel gravity with the function $F(T)$.
 The holographic dark energy model with the generalised entropy of the apparent horizon can be of interest for new cosmology.
-I Introduction
+1 Introduction
 To describe the current accelerating universe one can introduce in the Einstein equations the cosmological constant.
-It plays the role of dark energy leading to standard cosmology.
-Another way to describe acceleration of universe is to explore thermodynamics of apparent horizon Akbar ; Cai ; Cai0 ; Paranjape ; Sheykhi ; Cai1 ; Wang ; Jamil ; Gim ; Fan ; Agostino ; Sanchez because there is a correspondence between gravity and thermodynamics.
+It plays the role of dark energy leading to standard cosmology with its large scale homogeneity and isotropy.
+It is worth noting that our universe is homogeneous and isotropic on scales larger than $100$ Mpc [1].
+According to the Big Bang scenario there was homogeneous and isotropic distribution of a matter at high temperature and density about $15$ billion years ago.
+Then the universe has been expanding and cooling.
+The Friedmann equations with the positive cosmological constant lead to de Sitter spacetime which explains the current universe acceleration (due to the dark energy).
+Another way to describe acceleration of universe is to explore thermodynamics of apparent horizon [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] because there is a correspondence between gravity and thermodynamics.
 This is based on the fact that as was shown in Refs.
-Bekenstein ; Hawking that in black holes, entropy is connected with the horizon area and temperature is linked with the surface gravity Jacobson ; Padmanabhan ; Padmanabhan1 ; Hayward .
+[14, 15] that in black holes, the entropy is connected with the horizon area and temperature is linked with the surface gravity [16, 17, 18, 19].
 Also, Friedmann’s equations may be obtained with the help of the first law of apparent horizon thermodynamics.
-As a result, the apparent horizon of the FLRW space-time is represented as a thermodynamic system and it is a causal boundary Hayward ; Hayward1 ; Bak , and on this boundary the thermodynamics laws are satisfied Akbar ; Cai0 .
-In flat FLRW universe the apparent horizon is equal to the Hubble horizon.
-Different entropies Tsallis ; Ren ; Renyi ; Jahromi ; Barrow ; Kaniadakis ; Mejrhit ; Majhi were introduced, due to the long-range nature of gravity, that lead to the generalized Friedmann equations.
+As a result, the apparent horizon of the FLRW spacetime is represented as a thermodynamic system and it is a causal boundary [19, 20, 21], and on this boundary the thermodynamics laws are satisfied [2, 4].
+In spatially flat FLRW universe the apparent horizon is equal to the Hubble horizon.
+Different entropies [22, 23, 24, 25, 26, 27, 28] were introduced, due to the long-range nature of gravity, that lead to the generalized Friedmann equations.
+Thus, Tsallis entropy represents effective entropy for non-extensive thermodynamic systems with long range interactions.
+The Barrow entropy is similar to the Tsallis entropy and has an application in quantum gravity.
+The Rényi entropy is connected with the amount of information of the system.
+The Kaniadakis entropy appears in the consideration of relativistic statistical systems.
+The Sharma–Mittal entropy [27] is two-parameter entropy which is a combination of the Rényi and Tsallis entropies.
+The entropy of Loop Quantum Gravity [28] was used in black hole physics and cosmology.
+Previous studies of entropies and holographic dark energy models were considered in Refs.
+[29, 30, 31, 32, 33, 34, 35, 36].
 Motivated by success of this approach, it is of interest to consider other entropies to describe accelerating universe.
-Because entropies are sources of holographic energy densities they can describe the dark energy of the universe Pavon ; Landim .
-In this letter, we consider new apparent horizon entropy $S_{K}=S_{BH}/(1+\gamma S_{BH})$ where $S_{BH}$ is the Bekenstein–Hawking entropy and it is a non-extensive entropy measure.
+Because entropies are sources of holographic energy densities they can describe the dark energy of the universe [37, 38].
+In this paper, we consider new apparent horizon entropy $S_{K}=S_{BH}/(1+\gamma S_{BH})$ where $S_{BH}$ is the Bekenstein–Hawking entropy and it is a non-extensive entropy measure.
 When the Bekenstein–Hawking entropy becomes zero our entropy $S_{K}$ also vanishes, $S_{K}$ is the monotonically increasing function of $S_{BH}$ and is positive.
@@
 Here, we consider the case with equation of state (EoS) for barotropic perfect fluid, $p=w\rho$.
-We will obtain modified Friedmann’s equations with dynamical cosmological constant that leads to dark energy and the universe inflation.
-Our new holographic model with the non-additive entropy and non-extensive feature of space-time shows the possibility of the universe to accelerate in accordance with observations.
+We will obtain modified Friedmann’s equations with dynamical cosmological constant that leads to dark energy and the early and late time universe acceleration.
+Our new holographic model with the non-additive entropy shows the possibility of the universe to accelerate in accordance with observations.
 We assume units with $\hbar=c=k_{B}=1$.
-II New entropy
+2 New entropy
 We start with new entropy
-$S_{K}=-\sum_{i=1}^{W}\frac{p_{i}\ln{p_{i}}}{1-\gamma\ln{p_{i}}}),$ (1)
+$S_{K}=-\sum_{i=1}^{W}\frac{p_{i}\ln{p_{i}}}{1-\gamma\ln{p_{i}}},$ (1)
 where $W$ is a number of states.
@@
 In Eq.
-(2.1) the summation is performed over all possible system microstates.
-When $\gamma=0$ entropy (2.1) is converted into the Gibbs entropy
+(1) the summation is performed over all possible system microstates.
+When $\gamma=0$ entropy (1) is converted into the Gibbs entropy
 $S_{G}=-\sum_{i=1}^{W}p_{i}\ln(p_{i}).$ (2)
 If one assumes that each microstate is populated with equal probability, then $1/p_{i}=W$ ($i=1,2,...,W$) and Eq.
-(2.2) is converted into the Boltzmann entropy $S_{B}=\ln(W)$.
+(2) is converted into the Boltzmann entropy $S_{B}=\ln(W)$.
 Making use of $1/p_{i}=W$, we obtain from Eq.
-(2.1)
+(1)
 $S_{K}=\frac{\ln(W)}{1+\gamma\ln(W)}.$ (3)
-The Bekenstein–Hawking entropy $S_{BH}=\ln(W)$. Then from Eq. (2.3) one finds
+The Bekenstein–Hawking entropy is given by $S_{BH}=\ln(W)$. Then from Eq. (3) one finds
 $S_{K}=\frac{S_{BH}}{1+\gamma S_{BH}}.$ (4)
 Making use of Eq.
-(2.4), at $\gamma=0$, we obtain the Bekenstein–Hawking entropy $S_{BH}$.
+(4), at $\gamma=0$, we obtain the Bekenstein–Hawking entropy $S_{BH}$.
 If A and B are two probabilistically independent systems, we have $p^{A+B}_{ij}=p^{A}_{i}p^{B}_{j}$ and one finds nonadditive entropy, $S_{K}(A+B)\neq S_{K}(A)+S_{K}(B)$.
-III Apparent horizon thermodynamics
-Let us consider the FLRW flat universe which is described by the metric
+3 Apparent horizon thermodynamics
+Let us consider the FLRW spatially flat universe which is described by the metric
 $ds^{2}=-dt^{2}+a(t)^{2}(dr^{2}+r^{2}d\Omega_{2}^{2}).$ (5)
@@
 $R_{h}=\frac{1}{H},$ (6)
-where the Hubble parameter of the universe is $H=\dot{a}(t)/a(t)$, with dot over $a(t)$ being the derivative with respect to the cosmological time $t$.
+where the Hubble parameter of the universe, which measures the expansion rate, is $H=\dot{a}(t)/a(t)$, with dot over $a(t)$ being the derivative with respect to the cosmological time $t$.
 The first law of apparent horizon thermodynamics is formulated as
@@
 $E=\rho V_{h}=\frac{4\pi}{3}\rho R_{h}^{3}.$ (8)
-The change of the energy inside the apparent horizon is $-dE$. In Eq. (3.4) $\rho$ means the energy density of matter fields and the work density is Hayward ; Hayward1 ; Bak
+The change of the energy inside the apparent horizon is $dQ=-dE$. In Eq. (8) $\rho$ means the energy density of a matter and the work density is [19, 20, 21]
 $W=-\frac{1}{2}\mbox{Tr}(T^{\mu\nu})=\frac{1}{2}(\rho-p),$ (9)
@@
 $T_{h}=\frac{H}{2\pi}\left|1+\frac{\dot{H}}{2H^{2}}\right|.$ (10)
-Making use of first law of apparent horizon thermodynamics (3.3), and by combining Eqs. (3.4) and (3.5) with (3.6) , one finds
+Making use of first law of apparent horizon thermodynamics (7), and by combining Eqs. (8) and (9) with (10), one finds
 $\frac{H}{2\pi}\left|1+\frac{\dot{H}}{2H^{2}}\right|dS_{h}=-\frac{4\pi}{3H^{3}}d\rho+\frac{2\pi(\rho+p)}{H^{4}}dH.$ (11)
@@
 $\dot{\rho}=-3H(\rho+p),$ (12)
-and Eq. (3.7) we obtain
+and Eq. (11), we obtain
 $\frac{H}{2\pi}\left|1+\frac{\dot{H}}{2H^{2}}\right|\dot{S}_{h}=-\frac{4\pi\dot{\rho}}{3H^{3}}\left(1+\frac{\dot{H}}{2H^{2}}\right).$ (13)
-IV Generalized Friedmann’s equations
-Utilizing Eqs. (3.8), (3.9) and assuming that $01+\dot{H}/(2H^{2})>$ 0 we obtain
+4 Generalized Friedmann’s equations
+Utilizing Eqs. (12), (13) and assuming that $01+\dot{H}/(2H^{2})>$ 0 we obtain
 $\frac{H}{2\pi}\dot{S}_{h}=\frac{4\pi(\rho+p)}{H^{2}}.$ (14)
@@
 $S_{h}=S_{K}=\frac{S_{BH}}{1+\gamma S_{BH}},$ (15)
-where $S_{BH}=\pi R^{2}_{h}/G=\pi/(GH^{2})$, we obtain from Eq. (4.1) the generalized Friedmann equation
+where $S_{BH}=\pi R^{2}_{h}/G=\pi/(GH^{2})$, we obtain from Eq. (14) the generalized Friedmann equation
 $\frac{\dot{H}}{(1+\gamma\pi/(GH^{2}))^{2}}=-4\pi G(\rho+p).$ (16)
 At $\gamma=0$ in Eq.
-(4.3), we come to the first Friedmann equation for flat universe within Einstein gravity.
+(16), we come to the first Friedmann equation for spatial flat universe within Einstein gravity.
 Making use of Eq.
-(3.8) and after integration of Eq.
-(4.3), we find the second generalized Friedmann equation
+(12) and after integration of Eq.
+(16), we find the second generalized Friedmann equation
 $H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)=\frac{8\pi G}{3}\rho,$ (17)
@@
 At $\gamma=0$ ($b=0$) one obtains from Eq.
-(4.4) the Friedmann equation of general relativity.
+(17) the Friedmann equation of general relativity.
 We can represent Eq.
-(4.4) as follows:
-$H^{2}=\frac{8\pi G}{3}\rho+\frac{\Lambda_{eff}}{3}.$ (18)
+(17) as follows:
+$H^{2}=\frac{8\pi G}{3}\rho+\frac{\Lambda_{eff}}{3}$ (18)
 with
@@
 1 when $b$ increases, at fixed $H$, the dynamical cosmological constant $\Lambda_{eff}$ also increases and $\lim_{H\rightarrow 0}\Lambda_{eff}=3b$.
+Thus, at large apparent horizon radius $R_{h}$ (small $H$), we have the constant $\Lambda_{eff}$ explaining the current universe acceleration.
 By virtue of Eq.
-(4.5) one finds the density of dark energy
+(18) one finds the density of dark energy
 $\rho_{D}=\frac{3b}{8\pi G}\left[\frac{b}{H^{2}+b}+2\ln\left(\frac{H^{2}+b}{b}\right)\right].$ (20)
 Defining the normalized density parameters $\Omega_{m}=\rho/(3M_{P}^{2}H^{2})$ and $\Omega_{D}=\rho_{D}/(3M_{P}^{2}H^{2})$, where $M_{P}=1/\sqrt{8\pi G}$ is the reduced Plank mass, one obtains from Eqs.
-(4.4) and (4.7) that $\Omega_{m}+\Omega_{D}=1$.
+(17) and (20) that $\Omega_{m}+\Omega_{D}=1$.
 From Eq.
-(4.4) one obtains the normalized density for the matter $w=0$
-$\Omega_{m}=1-\frac{b^{2}}{(H^{2}H^{2}+b)}-2\frac{b}{H^{2}}\ln\left(\frac{H^{2}+b}{b}\right).$
+(17) one obtains the normalized density for the matter
+$\Omega_{m}=1-\frac{b^{2}}{H^{2}(H^{2}+b)}-2\frac{b}{H^{2}}\ln\left(\frac{H^{2}+b}{b}\right).$
 The plot of $\Omega_{m}$ versus $H$ is depicted in Fig. 2.
 Figure 2: The function $\Omega_{m}$ versus $H$ at $b==0.15,0.2,0.3$.
-Figure 2 shows that $\Omega_{m}$ increases as $b$ increases.
+Figure 2 shows that $\Omega_{m}$ increases as $b$ decreases.
 As $H\rightarrow\infty$ ($R_{h}\rightarrow 0$) we have $\Omega_{m}\rightarrow 1$.
 According to Fig.
-2 for the current era $\Omega_{m}=0.26$ that can be realized for some parameter $b$.
-We will imply here that there is no mutual interaction between various components of the cosmos, from ordinary conservation law, one obtains
+2 for the current era [39] $\Omega_{m0}=0.315$ that can be realized for some parameter $b$.
+We will imply here that there is no mutual interaction between various components of the cosmos, and from ordinary conservation law, one obtains
 $p_{D}=-\frac{\dot{\rho}_{D}}{3H}-\rho_{D}.$ (21)
-By virtue of Eqs. (4.7) and (4.8) we find the pressure corresponding to the dark energy
+By virtue of Eqs. (20) and (21) we find the pressure corresponding to the dark energy
 $p_{D}=-\frac{b(b+2H^{2})\dot{H}}{4\pi G(b+H^{2})^{2}}-$
 $\frac{3b}{8\pi G}\left[\frac{b}{H^{2}+b}+2\ln\left(\frac{H^{2}+b}{b}\right)\right].$ (22)
-Making use of Eqs, (4.3), (4.4) and (4.9) we obtain
+Making use of Eqs. (16), (17) and (22) we obtain
 $p_{D}=\frac{3b(b+2H^{2})(1+w)}{8\pi GH^{4}}\biggl{[}H^{2}-\frac{b}{H^{2}+b}-$
 $2\ln\left(\frac{H^{2}+b}{b}\right)\biggr{]}-\frac{3b}{8\pi G}\left[\frac{b}{H^{2}+b}+2\ln\left(\frac{H^{2}+b}{b}\right)\right].$ (23)
-With the help of Eqs. (4.7) and (4.10) one finds EoS for dark energy
+With the help of Eqs. (20) and (23) one finds EoS for dark energy
 $w_{D}=\frac{p_{D}}{\rho_{D}}=$
-$\frac{(b+2H^{2})(1+w)}{H^{4}}\left[\frac{H^{2}(H^{2}+b)}{b+2(H^{2}+b)\ln\left(\frac{H^{2}+b}{b}\right)}-1\right]-1.$ (24)
+$\frac{(b+2H^{2})(1+w)}{H^{4}}\left[\frac{H^{2}(H^{2}+b)}{b+2(H^{2}+b)\ln\left(\frac{H^{2}+b}{b}\right)}-b\right]-1.$ (24)
 The plot of $w_{D}$ versus $H$ is depicted in Fig. 3.
@@
 At large $H$ EoS parameter for dark energy $w_{D}$ approaches to $-1$.
-Right panel: In accordance with figure when parameter $b$ increases, at fixed $H$, EoS parameter for dark energy $w_{D}$ also increases and $\lim_{H\rightarrow\infty}w_{D}=-1$.
+Right panel: In accordance with figure when parameter $b$ increases, at fixed $H$, EoS parameter for dark energy $w_{D}$ decreases and $\lim_{H\rightarrow\infty}w_{D}=-1$.
 In accordance with Fig.
 3 (Left panel) at $b=1$ and $w=0,1/3,2/3$ when $w$ increases, at fixed $H$, EoS parameter for dark energy $w_{D}$ also increases.
-Figure 3 shows (Right panel) that at $w=1$ and $b=2,3,4$ if $b$ increases, at fixed $H$, $w_{D}$ aolso increases.
+Figure 3 shows (Right panel) that at $w=1$ and $b=2,3,4$ if $b$ increases, at fixed $H$, $w_{D}$ decreases.
 It follows from Eq.
-(4.11) that $\lim_{H\rightarrow\infty}w_{D}=-1$.
-Thus, dynamical cosmological constant leads to dark energy EoS $w_{D}=-1$ at large Habble parameter $H$.
+(24) that $\lim_{H\rightarrow\infty}w_{D}=-1$.
+Thus, dynamical cosmological constant leads to dark energy EoS $w_{D}=-1$ at large Hubble parameter $H$ (small $R_{h}$) which corresponds for the inflation era.
 The second law of thermodynamics requires that $\dot{S}_{K}\geq 0$.
 Then from Eq.
-(3.2) one obtains $\dot{S}_{BH}/(1+\gamma S_{BH})^{2}\geq 0$ and this leads to $\dot{S}_{BH}=-2\pi\dot{H}/(GH^{3})\geq 0$.
+(6) one obtains $\dot{S}_{BH}/(1+\gamma S_{BH})^{2}\geq 0$ and this leads to $\dot{S}_{BH}=-2\pi\dot{H}/(GH^{3})\geq 0$.
 As a result, we have the same requirement as for the Bekenstein–Hawking entropy.
@@
 Then from Eq.
-(3.3) we find $\rho+p\geq 0$ and for $\rho>0$ the EoS parameter $w\geq-1$.
+(7) we find $\rho+p\geq 0$ and for $\rho>0$ the EoS parameter $w\geq-1$.
 It is convenient to use the redshift $z=a_{0}/a(t)-1$, where $a_{0}$ corresponds to a scale factor at the current time.
-With the aid of continuity equation (3.8) and EoS $p=w\rho$ one obtains the matter density energy
+With the aid of continuity equation (12) and EoS $p=w\rho$ one obtains the matter density energy
 $\rho=\rho_{0}\left(\frac{1+z}{a_{0}}\right)^{3(1+w)},$ (25)
-with $\rho_{0}$ being the density energy of matter at the present time. From Eqs. (4.4) and (4.12) we find equation as follows:
+with $\rho_{0}$ being the density energy of matter at the present time. From Eqs. (20) and (25) we find equation as follows:
 $H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)=\frac{8\pi G\rho_{0}}{3}\left(\frac{1+z}{a_{0}}\right)^{3(1+w)}.$ (26)
@@
 At fixed $H$, when EoS parameter for the matter $w$ increases the redshift $z$ decreases.
-Right panel: In accordance with figure if parameter $b$ increases at fixed $z$ the Habble parameter also increases.
-When redshift $z$ increases the Habble parameter also increases.
+Right panel: In accordance with figure if parameter $b$ increases at fixed $z$ the Hubble parameter also increases.
+When redshift $z$ increases the Hubble parameter also increases.
 According to Fig.
 4 (Left panel) if parameter $w$ increases the redshift $z$ decreases.
-Figure 4 (Right panel) shows that when parameter $b$ increases at fixed z the Habble parameter also increases The deceleration parameter is given by
+Figure 4 (Right panel) shows that when parameter $b$ increases at fixed z the Hubble parameter also increases The deceleration parameter is given by
 $q=-\frac{\ddot{a}a}{\dot{a}^{2}}=-1-\frac{\dot{H}}{H^{2}}.$ (27)
-If $q<0$ we have the acceleration phase of the universe and when $q>0$ the universe decelerates. Making use of Eqs. (4.3), (4.12) and (4.14) we find
+If $q<0$ we have the acceleration phase of the universe and when $q>0$ the universe decelerates. Making use of Eqs. (16), (25) and (27) we find
 $q=\frac{4\pi G\rho_{0}(1+w)(H^{2}+b)^{2}}{H^{6}}\left(\frac{1+z}{a_{0}}\right)^{3(1+w)}-1.$ (28)
-Equation (4.15) defines the dependence of the deceleration parameter $q$ on redshift $z$. By virtue of Eqs. (4.13) and (4.15) we obtain the function of deceleration parameter $q$ on $H$
+Equation (28) defines the dependence of the deceleration parameter $q$ on redshift $z$. By virtue of Eqs. (26) and (28) we obtain the function of deceleration parameter $q$ on $H$
 $q=-1+$
 $\frac{3(1+w)(H^{2}+b)^{2}}{2H^{6}}\left(H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)\right).$ (29)
-We depicted the deceleration parameter $q$ versus the Hubble parameter $H$ for $\rho_{0}=a_{0}=1$ in Fig. 5.
+We depicted the deceleration parameter $q$ versus the Hubble parameter $H$ in Fig. 5.
 Figure 5: According to Fig.
-5 when deceleration parameter $q$ increases the Habble parameter $H$ also increases.
-Left panel: The deceleration parameter $q$ versus Habble parameter $H$ at $b=1$, $w=0,1/3$, 2/3, $\rho_{0}=a_{0}=1$.
+5 when deceleration parameter $q$ increases the Hubble parameter $H$ also increases.
+Left panel: The deceleration parameter $q$ versus Hubble parameter $H$ at $b=1$, $w=0,1/3$, 2/3, $\rho_{0}=a_{0}=1$.
 When EoS parameter for the matter $w$ increases, at fixed $b$ and $H$, the deceleration parameter $q$ also increases.
@@
 Here we have two phases, universe acceleration and deceleration.
-Figure 5 shows that there are two phases: universe acceleration and deceleration. Making use of Eq. (4.16) we find the asymptotic
+Figure 5 shows that there are two phases: universe acceleration and deceleration. Making use of Eq. (29) we find the asymptotic
 $\lim_{H\rightarrow\infty}q=\frac{3w+1}{2}.$ (30)
-Equation (4.17) shows that the asymptotic of the deceleration parameter as $H\rightarrow\infty$ ($R_{h}\rightarrow 0$) does not depend on the entropy parameter $\gamma$.
+Equation (30) shows that the asymptotic of the deceleration parameter as $H\rightarrow\infty$ ($R_{h}\rightarrow 0$) does not depend on the entropy parameter $\gamma$.
 It follows from Eq.
-(4.17) that when $w>-1/3$ ($q>0$) we have two phases: universe acceleration and deceleration.
-If $w<-1/3$ the eternal inflation takes place.
+(30) that when $w>-1/3$ ($q>0$), at small $R_{h}$, we have the universe deceleration.
+Thus, the inflation of the universe (the universe acceleration at small $R_{h}$) takes place at $w<-1/3$.
 By virtue of Eq.
-(4.13) we obtain the redshift $z$ as a function of Habble parameter
+(26) we obtain the redshift $z$ as a function of Hubble parameter
 $z=-1+$
@@
 From Eqs.
-(4.16) and (4.18) we found the approximate real and positive solutions for the transition redshifts $z_{t}$ (from acceleration to deceleration phases) at $q=0$, $G=1$, $w=0$ (for the matter dominated era) presented in Table 1.
-Table 1: The approximate solutions to Eqs. (3.16) and (3.18) for the transition redshifts $q=0$ at $G=1$, $a_{0}=\rho_{0}=1$, $w=0$.
+(29) and (31) we found the approximate real and positive solutions for the transition redshifts $z_{t}$ (from acceleration to deceleration phases) at $q=0$, $G=1$, $w=0$ (for the matter dominated era) presented in Table 1.
+Table 1: The approximate solutions to Eqs. (29) and (31) for the transition redshifts $q=0$ at $G=1$, $a_{0}=\rho_{0}=1$, $w=0$.
 $b$ 1.5 2 2.5 3 3.5 4 4.5 5 5.5 6 6.5
@@
 $z_{t}$ 0.07 0.18 0.27 0.35 0.42 0.49 0.55 0.60 0.65 0.70 0.75
-According to Table I when parameter $b$ increases (at fixed $w$) the Hubble parameter $H$ and reshift $z$ also increase for a point $q=0$ between universe acceleration and deceleration.
+According to Table I when parameter $b$ increases (at fixed $w$) the Hubble parameter $H$ and reshift $z_{t}$ also increase for a point $q=0$ between universe acceleration and deceleration.
 From Eqs.
-(4.16) and (4.18) for the current era $z=0$, we find solutions for the Habble parameter $H$ and deceleration parameter $q$ for varies $b$, given in Table II.
-Table 2: The solutions (approximate) to Eqs. (3.16) and (3.18) for the current era $z=0$ at $G=1$, $a_{0}=\rho_{0}=1$, $w=-2/3$.
+(29) and (31) for the current era $z=0$, we find solutions for the Hubble parameter $H$ and deceleration parameter $q$ for different $b$, given in Table II.
+Table 2: The solutions (approximate) to Eqs. (29) and (31) for the current era $z=0$ at $G=1$, $a_{0}=\rho_{0}=1$, $w=-2/3$.
 $b$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
@@
 In Table II the deceleration parameter $q$ is negative, and therefore, we have here the acceleration phase in the current time.
-The deceleration parameter at the current era is $q_{0}\approx-0.6$ Roos corresponding to parameter $b\approx 0.5$.
-By virtue of (3.16) we plotted the Habble parameter $H$ versus redshift $z$ in Fig.
+The deceleration parameter at the current era is [39] (see also [41]) $q_{0}\approx-0.535$ corresponding to parameter $b\approx 0.1$.
+Making use of Eqs.
+(29) and (31) we depicted the deceleration parameter $q$ versus redshift $z$ in Fig.
 6.
-Figure 6: Left panel: The Habble parameter $H$ versus $z$ at $b=0.1$, $w=-0.3,1/3$, 2/3, $G=1$, $\rho_{0}=a_{0}=1$.
-Figure 6 shows that when $z$ increases, $H$ also increases.
-When EoS parameter $w$ increases, at fixed $b$, the Habble parameter $H$ decreases.
-Right panel: When parameter $b$ increases, at fixed $w$, the Habble parameter $H$ also increases.
-In accordance with Fig.
-6, when redshift $z$ increases, the Habble parameter $H$ also increases.
-When EoS parameter $w$ increases, at fixed $b$, the Habble parameter $H$ also increases.
-Figure 6 shows that if parameter $b$ increases, at fixed $w$, the Habble parameter $H$ also increases.
-Making use of Eqs.
-(4.16) and (4.18) we depicted the deceleration parameter $q$ versus redshift $z$ in Fig.
-7.
 Figure 6: Left panel: The deceleration parameter $q$ versus $z$ at $b=0.1$, $w=-0.3,0$, 1/3, $G=1$, $\rho_{0}=a_{0}=1$.
-In accordance with Figure 7, when $z$ increases the deceleration parameter $q$ decreases.
+In accordance with Figure 6, when $z$ increases the deceleration parameter $q$ also increases.
 When EoS parameter for the matter $w$ increases, at fixed $b$, the deceleration parameter $q$ also increases.
@@
 Right panel: According to Fig.
-5 when parameter $b$ increases, at fixed $w$, the deceleration parameter $q$ also increases.
+6 when parameter $b$ increases, at fixed $w$ and $z$, the deceleration parameter $q$ decreases.
 We have here two phases: acceleration and deceleration for $b=0.4$ and $0.6$.
@@
 According to Fig.
-5 (Right panel), when parameter $b$ increases at fixed $w$ the deceleration parameter $q$ decreases.
+6 (Right panel), when parameter $b$ increases at fixed $w$ and $z$, the deceleration parameter $q$ decreases.
 One has here two phases, acceleration and deceleration for $b=0.4$ and $b=0.6$ but for $b=0.2$ we have the deceleration phase.
-V Conclusion
+4.1 Inflation
+Inflation corresponds to the accelerated expansion of the universe and gravity possesses a repulsive force, At the early stage of the universe, i.e.
+inflation, the matter is absent, $\rho=p=0$.
+The entropic energy (dark energy) density triggers the universe to accelerate.
+Then we obtain from Eq.
+(16) that $\dot{H}=0$, i.e.
+during inflation the Hubble parameter is a constant.
+As a result, the scale factor is given by $a(t)=a_{0}\exp(Ht)$ and corresponds to de Sitter spacetime.
+It should be noted that de Sitter stage corresponds to the eternal inflation.
+The Hubble parameter at the early universe (inflation) is $H\approx 10^{-3}M_{Pl}$ ($M_{Pl}=1/\sqrt{16\pi G}$).
+Then $GH^{2}\approx 2\cdot 10^{-8}$.
+From Eq.
+(17), at $\rho=0$ we obtain the equation as follows:
+$H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)=0.$ (32)
+Introducing new variable $x=H^{2}/b$, Eq. (32) becomes
+$x-\frac{1}{x+1}-2\ln(x+1)=0.$ (33)
+The non-trivial solution to Eq. (33) is given by $x=3.04002$ and we have the relation
+$H^{2}=3.04002b.$ (34)
+Knowing the Hubble parameter during universe inflation, $H\approx 10^{-3}M_{Pl}$, we can fix parameter $b$ from Eq.
+(34) and the entropy parameter $\gamma=bG/\pi$.
+Thus, we obtain $\gamma\approx 2\times 10^{-9}$.
+It is worth noting that to have a viable model of inflation one needs a smooth graceful exit into the deceleration stage [1].
+5 F(T)-gravity from generalized entropy
+In teleparallel gravity, a scalar torsion $T$ plays a role of a fundamental field similar to the curvature $R$ in Einstein’s general relativity theory [42, 43].
+To describe the inflationary era and the current universe accelerating expansion, one introduces a Lagrangian in the form of $F(T)$ in analogy with $F(R)$-gravity.
+The torsion scalar field $T$ is given by [44, 45]
+$T=S_{\rho}^{~{}\mu\nu}T^{\rho}_{~{}\mu\nu},$ (35)
+where tensors $S_{\rho}^{~{}\mu\nu}$ and $T^{\rho}_{~{}\mu\nu}$ are defined as follows:
+$S_{\rho}^{~{}\mu\nu}=\frac{1}{2}\left(K_{~{}~{}\rho}^{\mu\nu}+\delta^{\mu}_{\rho}T_{~{}~{}\alpha}^{\alpha\nu}-\delta^{\nu}_{\rho}T_{~{}~{}\alpha}^{\alpha\mu}\right),$
+$K^{\mu\nu}_{~{}~{}\rho}=-\frac{1}{2}\left(T^{\mu\nu}_{~{}~{}\rho}-T^{\nu\mu}_{~{}~{}\rho}-T^{~{}\mu\nu}_{\rho}\right),$
+$T^{\rho}_{~{}\mu\nu}=e^{\rho}_{i}\left(\partial_{\mu}e^{i}_{\nu}-\partial_{\nu}e^{i}_{\mu}\right),$ (36)
+where $e^{i}_{\nu}$ ($i=0,1,2,3$) is a vierbein field, so that the metric tensor is $g_{\mu\nu}=\eta_{ij}e^{i}_{\mu}e^{j}_{\nu}$ with $\eta_{ij}$ being the flat metric of the tangent spacetime.
+For FLRW metric (5), we have $e^{i}_{\mu}=\mbox{diag}(1,a,a,a)$.
+Then the torsion scalar field is given by $T=-6H^{2}$.
+From the action by variation with respect to $e^{i}_{\mu}$, where the Lagrangian is $F(T)$, one obtains [46]
+$\frac{1}{6}\left[F(T)-2TF^{\prime}(T)\right]|_{T=-6H^{2}}=\left(\frac{8\pi G}{3}\right)\rho.$ (37)
+Making use of equations (17) and (37) we find
+$\left[F(T)-2TF^{\prime}(T)\right]=-T-\frac{36b^{2}}{6b-T}-12b\ln\left(1-\frac{T^{2}}{6b}\right).$ (38)
+By integration of Eq. (38) we obtain the function $F(T)$:
+$F(T)=T-3\sqrt{6bT}~{}\mbox{tanh}^{-1}\sqrt{\frac{T}{6b}}$
+$-12b\ln\left(1-\frac{T}{6b}\right)+C\sqrt{T}-6b,$ (39)
+where $C$ is the integration constant. Thus, the teleparallel gravity with the function (39) corresponds to entropic cosmology with entropy (15) under consideration.
+6 Conclusion
 Considering a novel entropy $S_{K}=S_{BH}/(1+\gamma S_{BH})$ we have a property similar to the Bekenstein–Hawking entropy $S_{BH}$.
@@
 The $S_{K}$ vanishes when the apparent horizon radius $R_{h}$ is zero and it increases monotonically when the apparent horizon radius $R_{h}$ increases.
-The barotropic perfect fluid and flat FLRW universe were studied.
+The barotropic perfect fluid and spatial flat FLRW universe were studied.
 We have obtained, from first law of apparent horizon thermodynamics, the modified Friedmann’s equations.
@@
 The EoS parameter for dark energy $w_{D}=p_{D}/\rho_{D}$ was calculated and it was shown that $\lim_{H\rightarrow\infty}w_{D}=-1$.
-The universe can have two phases, acceleration and deceleration, that is due to holographic dark energy.
-It should be noted that Barrow and Tsallis entropies also lead to cosmology which ia due to Einsten’s equations with the dynamical cosmological constant Gennaro .
-We computed the deceleration parameter that shows the possibility, for some model parameters, to describe inflation and acceleration at current era.
+Thus, at the small apparent horizon radius $R_{h}$ the de Sitter stage is realized which describes the inflation.
+In the model under consideration the universe can have two phases, acceleration and deceleration, that is due to holographic dark energy.
+It should be noted that Barrow and Tsallis entropies also lead to cosmology which ia due to Einstein’s equations with the dynamical cosmological constant [40].
+We computed the deceleration parameter that shows the possibility, for some model parameters, to describe the acceleration at the current era.
 The transition redshifts (when $q=0$) was calculated (given in Table I) for some parameters.
-We show (see Table II) that at $b\approx 0.5$ and $w=-2/3$ we have the deceleration parameter $q_{0}\approx-0.6$ which was observed in universe at the current era Roos .
-Thus, our approach, based on new entropy and leading to modified Friedmann equations, can description universe inflation and late time universe acceleration.
-It is worth mentioning that the results obtained here are similar to those in another model with different entropy Kruglov .
-Data availability statement: No data was used for the research described in the article.
+We show (see Table II) that at $b\approx 0.1$ and $w=-2/3$ we have the deceleration parameter $q_{0}\approx-0.535$ which was observed in universe at the current era [39].
+Thus, our approach, based on new entropy and leading to modified Friedmann equations, can describe universe inflation and late time universe acceleration.
+It has been proven that entropic cosmology with our entropy proposed is equivalent to cosmology based on the teleparallel gravity with the function $F(T)$.

### v4

@@
 Here, we consider new nonadditive entropy of the apparent horizon $S_{K}=S_{BH}/(1+\gamma S_{BH})$ with $S_{BH}$ being the Bekenstein–Hawking entropy.
-This is an alternative of the Rényi and Tsallis entropies, that allow us by utilising the holographic principle to develop of a new model of entropic (holographic) dark energy.
+This is an alternative of the Rényi and Tsallis entropies, that allows us, by utilising the holographic principle, to develop a new model of entropic (holographic) dark energy.
 When $\gamma\rightarrow 0$ our entropy becomes the Bekenstein–Hawking entropy $S_{BH}$.
-The generalised Friedmann’s equations for Friedmann–Lemaître–Robertson–Walker (FLRW) space-time for the barotropic matter fluid with $p=w\rho$ were obtained.
-We compute the dark energy pressure $p_{D}$, density energy $\rho_{D}$ and the deceleration parameter corresponding to our model.
-From the second modified Friedmann’s equation a dynamical cosmological constant was obtained.
-We show that at some model parameters $w$ and $\gamma$ there are two phases, universe acceleration and deceleration or the phase of the eternal inflation.
-Thus, our model, by virtue of the holographic principle, can describe the universe inflation and late time of the universe acceleration.
-It is shown that entropic cosmology with our entropy proposed is equivalent to cosmology based on the teleparallel gravity with the function $F(T)$.
+The generalised Friedmann equations for Friedmann–Lemaître–Robertson–Walker (FLRW) spacetime for the barotropic matter fluid with $p=w\rho$ were obtained.
+We compute the dark energy pressure $p_{D}$, density of energy $\rho_{D}$, the normalized density parameters $\Omega_{D}$, $\Omega_{m}$ and the deceleration parameter $q$ of the universe corresponding to our model.
+From the second modified Friedmann equation a dynamical cosmological constant was obtained.
+We show that at some model parameters $w$ and $\gamma$ we obtain $\Omega_{m0}\approx 0.315$ and $q_{0}\approx-0.535$ which are in agreement with the Planck data.
+[40].
+It was shown that the model under consideration possesses the phantom divide for the EoS of dark energy.
+Thus, our model, by virtue of the holographic principle, can describe the universe inflation and the late time of the universe acceleration.
+It is shown that entropic cosmology with our entropy proposed is equivalent to cosmology based on the teleparallel gravity with the function $F(T)$ obtained.
 The holographic dark energy model with the generalised entropy of the apparent horizon can be of interest for new cosmology.
 1 Introduction
-To describe the current accelerating universe one can introduce in the Einstein equations the cosmological constant.
+To describe the current universe acceleration, one can introduce in the Einstein equation the cosmological constant.
 It plays the role of dark energy leading to standard cosmology with its large scale homogeneity and isotropy.
@@
 This is based on the fact that as was shown in Refs.
-[14, 15] that in black holes, the entropy is connected with the horizon area and temperature is linked with the surface gravity [16, 17, 18, 19].
+[14, 15], in black holes the entropy is connected with the horizon area and the temperature is linked with the surface gravity [16, 17, 18, 19].
 Also, Friedmann’s equations may be obtained with the help of the first law of apparent horizon thermodynamics.
@@
 In spatially flat FLRW universe the apparent horizon is equal to the Hubble horizon.
-Different entropies [22, 23, 24, 25, 26, 27, 28] were introduced, due to the long-range nature of gravity, that lead to the generalized Friedmann equations.
+Different entropies [22, 23, 24, 25, 26, 27, 28, 29] were introduced, due to the long-range nature of gravity, that lead to the generalized Friedmann equations.
 Thus, Tsallis entropy represents effective entropy for non-extensive thermodynamic systems with long range interactions.
@@
 The entropy of Loop Quantum Gravity [27] was used in black hole physics and cosmology.
+The entropies proposed in Ref.
+[28, 29] are used in entropic cosmology and does not possesses a singularity at the Hubble rate $H=0$.
 Previous studies of entropies and holographic dark energy models were considered in Refs.
@@
 We will obtain modified Friedmann’s equations with dynamical cosmological constant that leads to dark energy and the early and late time universe acceleration.
+The model under consideration is a modification of the model [29] and uses dimensionless variables that is convenient for the analyse.
 Our new holographic model with the non-additive entropy shows the possibility of the universe to accelerate in accordance with observations.
@@
 $S_{G}=-\sum_{i=1}^{W}p_{i}\ln(p_{i}).$ (2)
-If one assumes that each microstate is populated with equal probability, then $1/p_{i}=W$ ($i=1,2,...,W$) and Eq.
+If one assumes that each microstate is populated with equal probability, then $1/p_{i}=W$ (i=1,2,,Wi=1,2,...,W) and Eq.
 (2) is converted into the Boltzmann entropy $S_{B}=\ln(W)$.
 Making use of $1/p_{i}=W$, we obtain from Eq.
-(1)
+(1) equation as follows:
 $S_{K}=\frac{\ln(W)}{1+\gamma\ln(W)}.$ (3)
@@
 Making use of Eq.
-(4), at $\gamma=0$, we obtain the Bekenstein–Hawking entropy $S_{BH}$.
-If A and B are two probabilistically independent systems, we have $p^{A+B}_{ij}=p^{A}_{i}p^{B}_{j}$ and one finds nonadditive entropy, $S_{K}(A+B)\neq S_{K}(A)+S_{K}(B)$.
+(4), at $\gamma=0$, we obtain the Bekenstein–Hawking entropy $S_{BH}=A/(4G)$, where $A=4\pi R_{h}^{2}$ is the area of the horizon.
+If $B$ and $C$ are two probabilistically independent systems, we have $p^{B+C}_{ij}=p^{B}_{i}p^{C}_{j}$ and one finds nonadditive entropy, $S_{K}(B+C)\neq S_{K}(B)+S_{K}(C)$.
+It is worth noting that the Bekenstein–Hawking entropy $S_{BH}$ as well as entropies [22, 23, 24, 25] possess a singularity when the Hubble parameter vanishes, $H=0$, but our entropy becomes $S_{K}=1/\gamma$ so that the singularity is absent.
 3 Apparent horizon thermodynamics
@@
 $E=\rho V_{h}=\frac{4\pi}{3}\rho R_{h}^{3}.$ (8)
-The change of the energy inside the apparent horizon is $dQ=-dE$. In Eq. (8) $\rho$ means the energy density of a matter and the work density is [19, 20, 21]
+The change of the energy inside the apparent horizon is dQ=dEdQ=-dE. In Eq. (8) $\rho$ means the energy density of a matter and the work density is [19, 20, 21]
 $W=-\frac{1}{2}\mbox{Tr}(T^{\mu\nu})=\frac{1}{2}(\rho-p),$ (9)
@@
 4 Generalized Friedmann’s equations
-Utilizing Eqs. (12), (13) and assuming that $01+\dot{H}/(2H^{2})>$ 0 we obtain
+Utilizing Eqs. (12), (13) and assuming that $1+\dot{H}/(2H^{2})>0$ we obtain
 $\frac{H}{2\pi}\dot{S}_{h}=\frac{4\pi(\rho+p)}{H^{2}}.$ (14)
@@
 At $\gamma=0$ in Eq.
-(16), we come to the first Friedmann equation for spatial flat universe within Einstein gravity.
+(16), we come to the first Friedmann equation for spatial flat universe within Einstein’s gravity.
 Making use of Eq.
@@
 $\Lambda_{eff}=\frac{3b^{2}}{H^{2}+b}+6b\ln\left(\frac{H^{2}+b}{b}\right).$ (19)
-The $\Lambda_{eff}$ represents the dynamical (effective) cosmological constant. We depicted the dynamical cosmological constant $\Lambda_{eff}$ versus $H$ at $b=1,2,3$ in Fig. 1.
+The $\Lambda_{eff}$ represents the dynamical (effective) cosmological constant.
+We depicted the dynamical cosmological constant $\Lambda_{eff}$ versus $H$ at $b=1,2,3$ in Fig.
+1.
+We use here the Planckian units with $G=c=\hbar=k_{B}=1$ [1].
 Figure 1: The function $\Lambda_{eff}$ versus $H$ at $b=\pi\gamma/G=1,2,3$.
@@
 $\rho_{D}=\frac{3b}{8\pi G}\left[\frac{b}{H^{2}+b}+2\ln\left(\frac{H^{2}+b}{b}\right)\right].$ (20)
-Defining the normalized density parameters $\Omega_{m}=\rho/(3M_{P}^{2}H^{2})$ and $\Omega_{D}=\rho_{D}/(3M_{P}^{2}H^{2})$, where $M_{P}=1/\sqrt{8\pi G}$ is the reduced Plank mass, one obtains from Eqs.
+Defining the normalized density parameters $\Omega_{m}=\rho/(3M_{P}^{2}H^{2})$ and $\Omega_{D}=\rho_{D}/(3M_{P}^{2}H^{2})$, where $M_{P}=1/\sqrt{8\pi G}$ is the reduced Planck mass, one obtains from Eqs.
 (17) and (20) that $\Omega_{m}+\Omega_{D}=1$.
 From Eq.
-(17) one obtains the normalized density for the matter
-$\Omega_{m}=1-\frac{b^{2}}{H^{2}(H^{2}+b)}-2\frac{b}{H^{2}}\ln\left(\frac{H^{2}+b}{b}\right).$
-The plot of $\Omega_{m}$ versus $H$ is depicted in Fig. 2.
-Figure 2: The function $\Omega_{m}$ versus $H$ at $b==0.15,0.2,0.3$.
-Figure 2 shows that $\Omega_{m}$ increases as $b$ decreases.
-As $H\rightarrow\infty$ ($R_{h}\rightarrow 0$) we have $\Omega_{m}\rightarrow 1$.
-According to Fig.
-2 for the current era [39] $\Omega_{m0}=0.315$ that can be realized for some parameter $b$.
+(17) we obtain the normalized density for the matter
+$\Omega_{m}=1-\frac{b^{2}}{H^{2}(H^{2}+b)}-2\frac{b}{H^{2}}\ln\left(\frac{H^{2}+b}{b}\right).$ (21)
+By introducing new dimensionless variable $x=H^{2}/b$, Eq. (21) is rewritten as
+$\Omega_{m}=1-\frac{1}{x(1+x)}-\frac{2}{x}\ln(1+x).$ (22)
+The plot of $\Omega_{m}$ versus $x$ is depicted in Fig. 2.
+Figure 2: The function $\Omega_{m}$ versus $x$. Figure 2 shows that $\Omega_{m}$ increases when $x$ increases. As $x\rightarrow\infty$ ($R_{h}\rightarrow 0$) we have $\Omega_{m}\rightarrow 1$.
+In accordance with Fig.
+2 as $x\rightarrow\infty$ ($H\rightarrow\infty$, $R_{h}\rightarrow 0$) one has $\Omega_{m}\rightarrow 1$.
+According to the Planck data [40], for the current era, $\Omega_{m0}\approx 0.315$.
+The solution to Eq.
+(22) at $\Omega_{m0}=0.315$ is given by $x\approx 5.819$.
+Then we obtain the entropy parameter $\gamma$,
+$\gamma=\frac{bG}{\pi}=\frac{GH_{0}^{2}}{5.819\pi}\approx 0.055~GH_{0}^{2}.$ (23)
+The current value of the Hubble rate, according to the Planck data [40], is $H_{0}\approx 67.4$ km/s/Mpc.
 We will imply here that there is no mutual interaction between various components of the cosmos, and from ordinary conservation law, one obtains
 $p_{D}=-\frac{\dot{\rho}_{D}}{3H}-\rho_{D}.$ (24)
-By virtue of Eqs. (20) and (21) we find the pressure corresponding to the dark energy
-$p_{D}=-\frac{b(b+2H^{2})\dot{H}}{4\pi G(b+H^{2})^{2}}-$
-$\frac{3b}{8\pi G}\left[\frac{b}{H^{2}+b}+2\ln\left(\frac{H^{2}+b}{b}\right)\right].$ (22)
+By virtue of Eqs. (2o) and (24) we find the pressure corresponding to the dark energy
+$p_{D}=-\frac{b(b+2H^{2})\dot{H}}{4\pi G(b+H^{2})^{2}}-\frac{3b}{8\pi G}\left[\frac{b}{H^{2}+b}+2\ln\left(\frac{H^{2}+b}{b}\right)\right].$ (25)
 Making use of Eqs. (16), (17) and (25) we obtain
@@
 With the help of Eqs. (20) and (26) one finds EoS for dark energy
-$w_{D}=\frac{p_{D}}{\rho_{D}}=$
-$\frac{(b+2H^{2})(1+w)}{H^{4}}\left[\frac{H^{2}(H^{2}+b)}{b+2(H^{2}+b)\ln\left(\frac{H^{2}+b}{b}\right)}-b\right]-1.$ (24)
-The plot of $w_{D}$ versus $H$ is depicted in Fig. 3.
-Figure 3: Left panel: The function $w_{D}$ versus $H$ at $b=1$, $w=0,1/3,2/3$, $G=1$, $\rho_{0}=a_{0}=1$.
-According to Fig.
-3 the $w_{D}$ increases as EoS parameter for the matter $w$ increases, at fixed $H$.
-At large $H$ EoS parameter for dark energy $w_{D}$ approaches to $-1$.
-Right panel: In accordance with figure when parameter $b$ increases, at fixed $H$, EoS parameter for dark energy $w_{D}$ decreases and $\lim_{H\rightarrow\infty}w_{D}=-1$.
-In accordance with Fig.
-3 (Left panel) at $b=1$ and $w=0,1/3,2/3$ when $w$ increases, at fixed $H$, EoS parameter for dark energy $w_{D}$ also increases.
-Figure 3 shows (Right panel) that at $w=1$ and $b=2,3,4$ if $b$ increases, at fixed $H$, $w_{D}$ decreases.
+$w_{D}=\frac{p_{D}}{\rho_{D}}=\frac{(b+2H^{2})(1+w)}{H^{4}}\left[\frac{H^{2}(H^{2}+b)}{b+2(H^{2}+b)\ln\left(\frac{H^{2}+b}{b}\right)}-b\right]-1.$ (27)
+Making use of the variable $x=H^{2}/b$ we represent Eq. (27) as follows:
+$w_{D}=\frac{p_{D}}{\rho_{D}}=\frac{(1+2x)(1+w)}{x}\left[\frac{x+1}{1+2(x+1)\ln(x+1)}-\frac{1}{x}\right]-1.$ (28)
+The plot of $w_{D}$ versus $x$ is depicted in Fig. 3.
+Figure 3: The function $w_{D}$ versus $x=H^{2}/b$ at $w=0,1/3,2/3$. At large $x$ the EoS parameter for dark energy $w_{D}$ approaches to 1-1, $\lim_{x\rightarrow\infty}w_{D}=-1$.
 It follows from Eq.
-(24) that $\lim_{H\rightarrow\infty}w_{D}=-1$.
+(28) that $\lim_{x\rightarrow\infty}w_{D}=-1$.
 Thus, dynamical cosmological constant leads to dark energy EoS $w_{D}=-1$ at large Hubble parameter $H$ (small $R_{h}$) which corresponds for the inflation era.
+Figure 3 shows that there is the phantom phase with $w_{D}<-1$.
+From Eq.
+(28) we obtain the equation for the phantom divide $w_{D}=-1$ as follows:
+$x(1+x)-2(x+1)\ln(x+1)-1=0.$ (29)
+The solution to Eq.
+(29) is $x=H^{2}/b\approx 3.04$ (see Fig.
+3).
+The phantom divide does not depend on the EoS of the matter $w$.
+Thus, our model realizes the crossing of the phantom divide for the EoS of the dark energy.
 The second law of thermodynamics requires that $\dot{S}_{K}\geq 0$.
 Then from Eq.
-(6) one obtains $\dot{S}_{BH}/(1+\gamma S_{BH})^{2}\geq 0$ and this leads to $\dot{S}_{BH}=-2\pi\dot{H}/(GH^{3})\geq 0$.
+(6) one finds $\dot{S}_{BH}/(1+\gamma S_{BH})^{2}\geq 0$ and this leads to $\dot{S}_{BH}=-2\pi\dot{H}/(GH^{3})\geq 0$.
 As a result, we have the same requirement as for the Bekenstein–Hawking entropy.
@@
 It is convenient to use the redshift $z=a_{0}/a(t)-1$, where $a_{0}$ corresponds to a scale factor at the current time.
-With the aid of continuity equation (12) and EoS $p=w\rho$ one obtains the matter density energy
-$\rho=\rho_{0}\left(\frac{1+z}{a_{0}}\right)^{3(1+w)},$ (25)
+With the aid of continuity equation (12) and EoS $p=w\rho$, one obtains the matter density energy
+$\rho=\rho_{0}(1+z)^{3(1+w)},$ (30)
 with $\rho_{0}$ being the density energy of matter at the present time. From Eqs. (17) and (30) we find equation as follows:
-$H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)=\frac{8\pi G\rho_{0}}{3}\left(\frac{1+z}{a_{0}}\right)^{3(1+w)}.$ (26)
-The Hubble parameter versus redshift $z$ for $G=1$, $\rho_{0}=a_{0}=1$ is plotted in Fig. 4.
-Figure 4: Left panel: The function $H$ versus $z$ at $b=0.1$, $w=1/3,0,-0.1$, $G=1$, $\rho_{0}=a_{0}=1$.
+$H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)=\frac{8\pi G\rho_{0}}{3}(1+z)^{3(1+w)}.$ (31)
+From Eq. (31) we obtain the redshift
+$z=\left[\frac{3}{8\pi\rho_{0}G}\left(H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)\right)\right]^{1/(3(1+w))}-1.$ (32)
+By introducing dimensionless parameters $\bar{H}=H/\sqrt{G\rho_{0}}$, $\bar{b}=b/(G\rho_{0})$, Eq. (32) becomes
+$z=\left[\frac{3}{8\pi}\left(\bar{H}^{2}-\frac{\bar{b}^{2}}{\bar{H}^{2}+\bar{b}}-2\bar{b}\ln\left(\frac{\bar{H}^{2}+\bar{b}}{\bar{b}}\right)\right)\right]^{1/(3(1+w))}-1.$ (33)
+The reduced Hubble parameter $\bar{H}$ versus redshift $z$ is plotted in Fig. 4.
+Figure 4: Left panel: The function $\bar{H}$ versus $z$ at $\bar{b}=0.1$, w=1/3,0,0.1w=1/3,0,-0.1.
 According to Fig.
-4 $H$ increases as $z$ increases.
-At fixed $H$, when EoS parameter for the matter $w$ increases the redshift $z$ decreases.
-Right panel: In accordance with figure if parameter $b$ increases at fixed $z$ the Hubble parameter also increases.
-When redshift $z$ increases the Hubble parameter also increases.
+4, $\bar{H}$ increases as $z$ increases.
+At fixed $\bar{H}$, when EoS parameter for the matter $w$ increases the redshift $z$ decreases.
+Right panel: In accordance with figure if parameter $\bar{b}$ increases, at fixed $z$, the reduced Hubble parameter $\bar{H}$ also increases.
+When redshift $z$ increases the reduced Hubble parameter also increases.
 According to Fig.
-4 (Left panel) if parameter $w$ increases the redshift $z$ decreases.
-Figure 4 (Right panel) shows that when parameter $b$ increases at fixed z the Hubble parameter also increases The deceleration parameter is given by
+4 (Left panel) if parameter $w$ increases, at fixed $\bar{H}$, the redshift $z$ decreases.
+Figure 4 (Right panel) shows that when parameter $\bar{b}$ increases, at fixed z, the reduced Hubble parameter $\bar{H}$ also increases.
+The deceleration parameter is given by
 $q=-\frac{\ddot{a}a}{\dot{a}^{2}}=-1-\frac{\dot{H}}{H^{2}}.$ (34)
 If $q<0$ we have the acceleration phase of the universe and when $q>0$ the universe decelerates. Making use of Eqs. (16), (30) and (34) we find
-$q=\frac{4\pi G\rho_{0}(1+w)(H^{2}+b)^{2}}{H^{6}}\left(\frac{1+z}{a_{0}}\right)^{3(1+w)}-1.$ (28)
+$q=\frac{4\pi G\rho_{0}(1+w)(H^{2}+b)^{2}}{H^{6}}\left(1+z\right)^{3(1+w)}-1.$ (35)
 Equation (35) defines the dependence of the deceleration parameter $q$ on redshift $z$. By virtue of Eqs. (31) and (35) we obtain the function of deceleration parameter $q$ on $H$
-$q=-1+$
-$\frac{3(1+w)(H^{2}+b)^{2}}{2H^{6}}\left(H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)\right).$ (29)
-We depicted the deceleration parameter $q$ versus the Hubble parameter $H$ in Fig. 5.
-Figure 5: According to Fig.
-5 when deceleration parameter $q$ increases the Hubble parameter $H$ also increases.
-Left panel: The deceleration parameter $q$ versus Hubble parameter $H$ at $b=1$, $w=0,1/3$, 2/3, $\rho_{0}=a_{0}=1$.
-When EoS parameter for the matter $w$ increases, at fixed $b$ and $H$, the deceleration parameter $q$ also increases.
+$q=\frac{3(1+w)(H^{2}+b)^{2}}{2H^{6}}\left(H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)\right)-1.$ (36)
+Making use of dimensionless variable $x=H^{2}/b$, Eq. (36) becomes
+$q=\frac{3(1+w)(1+x)^{2}}{2x^{2}}\left(1-\frac{1}{x(1+x)}-\frac{2}{x}\ln(1+x)\right)-1.$ (37)
+Taking into account $x=H_{0}^{2}/b\approx 5.819$ ($\gamma\approx H_{0}^{2}G/(5.819\pi)$) which gives the normalized density for the matter field at the current time $\Omega_{m0}\approx 0.315$ and the deceleration parameter $q_{0}\approx-0.535$ [40], we obtain the solution to Eq.
+(37) for the EoS parameter of the matter $w\approx-0.2833$.
+Thus, the current values of $\Omega_{m0}$ and $q_{0}$ are in accordance with the Planck data at the parameter $b=H_{0}^{2}/5.819$, where $H_{0}\approx 67$ km/s/Mpc and the EoS parameter of the matter $w\approx-0.2833$.
+We depicted the deceleration parameter $q$ versus the $x$ in Fig. 5.
+Figure 5: The deceleration parameter $q$ versus $x$ at $w=0,1/3,2/3$.
+According to figure the deceleration parameter $q$ increases when the parameter $x$ increases.
+When the EoS parameter of the matter $w$ increases, at fixed $x$, the deceleration parameter $q$ also increases.
 There are two phases: universe acceleration $q<0$ and deceleration $q>0$.
-Right panel: in accordance with figure, when parameter $b$ increases, at fixed $w$ and $H$, the deceleration parameter $q$ decreases.
-Here we have two phases, universe acceleration and deceleration.
-Figure 5 shows that there are two phases: universe acceleration and deceleration. Making use of Eq. (29) we find the asymptotic
+Figure 5 shows that there are two phases: the universe acceleration and deceleration.
+Making use of Eq. (36) we find the asymptotic
 $\lim_{H\rightarrow\infty}q=\frac{3w+1}{2}.$ (38)
@@
 It follows from Eq.
-(30) that when $w>-1/3$ ($q>0$), at small $R_{h}$, we have the universe deceleration.
-Thus, the inflation of the universe (the universe acceleration at small $R_{h}$) takes place at $w<-1/3$.
+(38) that when w>1/3w>-1/3 ($q>0$), at small $R_{h}$, we have the universe deceleration.
+Thus, the inflation of the universe (the universe acceleration at small $R_{h}$) takes place at w<1/3w<-1/3.
 By virtue of Eq.
-(26) we obtain the redshift $z$ as a function of Hubble parameter
-$z=-1+$
-$a_{0}\left[\frac{3}{8\pi\rho_{0}G}\left(H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)\right)\right]^{1/(3(1+w))}.$ (31)
-From Eqs.
-(29) and (31) we found the approximate real and positive solutions for the transition redshifts $z_{t}$ (from acceleration to deceleration phases) at $q=0$, $G=1$, $w=0$ (for the matter dominated era) presented in Table 1.
-Table 1: The approximate solutions to Eqs. (29) and (31) for the transition redshifts $q=0$ at $G=1$, $a_{0}=\rho_{0}=1$, $w=0$.
-$b$ 1.5 2 2.5 3 3.5 4 4.5 5 5.5 6 6.5
-$H$ 4.26 4.92 5.50 6.03 6.51 6.96 7.38 7.78 8.16 8.53 8.88
-$z_{t}$ 0.07 0.18 0.27 0.35 0.42 0.49 0.55 0.60 0.65 0.70 0.75
-According to Table I when parameter $b$ increases (at fixed $w$) the Hubble parameter $H$ and reshift $z_{t}$ also increase for a point $q=0$ between universe acceleration and deceleration.
-From Eqs.
-(29) and (31) for the current era $z=0$, we find solutions for the Hubble parameter $H$ and deceleration parameter $q$ for different $b$, given in Table II.
-Table 2: The solutions (approximate) to Eqs. (29) and (31) for the current era $z=0$ at $G=1$, $a_{0}=\rho_{0}=1$, $w=-2/3$.
-$b$ 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
-$H$ 3.05 3.15 3.25 3.33 3.40 3.47 3.54 3.60 3.66
-$q$0.54 -0.56 -0.58 -0.59 -0.61 -0.62 -0.63 -0.64 -0.65
-In Table II the deceleration parameter $q$ is negative, and therefore, we have here the acceleration phase in the current time.
-The deceleration parameter at the current era is [39] (see also [41]) $q_{0}\approx-0.535$ corresponding to parameter $b\approx 0.1$.
-Making use of Eqs.
-(29) and (31) we depicted the deceleration parameter $q$ versus redshift $z$ in Fig.
-6.
-Figure 6: Left panel: The deceleration parameter $q$ versus $z$ at $b=0.1$, $w=-0.3,0$, 1/3, $G=1$, $\rho_{0}=a_{0}=1$.
-In accordance with Figure 6, when $z$ increases the deceleration parameter $q$ also increases.
-When EoS parameter for the matter $w$ increases, at fixed $b$, the deceleration parameter $q$ also increases.
-At $w=-0.3$ there is two phases, universe acceleration $q<0$ and deceleration $q>0$ and at $w=0$ and $1/3$ we have only the deceleration phase.
-Right panel: According to Fig.
-6 when parameter $b$ increases, at fixed $w$ and $z$, the deceleration parameter $q$ decreases.
-We have here two phases: acceleration and deceleration for $b=0.4$ and $0.6$.
-For $b=0.2$ one has only the deceleration phase.
-Figure 6 shows that when $z$ increases $q$ also increases.
-When EoS parameter for the matter $w$ increases, at fixed $b$ (Left panel), the deceleration parameter $q$ also increases.
-At $w=-0.3$ there are two phases, universe acceleration $q<0$ and deceleration $q>0$ but at $w=0$ and $w=1/3$ one has only the deceleration phase.
+(37) at $q=0$ we obtain the equation for the transition phase
+$w=\frac{2x^{3}}{3(1+x)\left(x(1+x)-2(1+x)\ln(1+x)-1\right)}-1.$ (39)
+We plotted the EoS parameter for the matter $w$ versus $x$ in Fig. 6.
+Figure 6: The EoS parameter for the matter $w$ verses $x$ at $q=0$.
+When $x$ increases from $x=0$ ($H=0$, the EoS parameter $w$ decreases from w=1w=-1 (de Sitter space) and becomes the fantom space.
+Thus, the transition from the acceleration phase ar large $R_{h}$ (small $H$) into the deceleration phase is possible only for w<1w<-1.
+Figure 6 shows that when $x=H^{2}/b$ increases from $x=0$ the EoS parameter $w$ decreases from w=1w=-1 (de Sitter space) and becomes the phantom space.
+As a result, the transition from the acceleration phase ar large $R_{h}$ (small $H$) into the deceleration phase is possible only for fantom space w<1w<-1.
 According to Fig.
-6 (Right panel), when parameter $b$ increases at fixed $w$ and $z$, the deceleration parameter $q$ decreases.
-One has here two phases, acceleration and deceleration for $b=0.4$ and $b=0.6$ but for $b=0.2$ we have the deceleration phase.
+5, for the bigger $x$ ($x>3$) the transition from $q<0$ to $q>0$ takes place at $w>0$.
 4.1 Inflation
-Inflation corresponds to the accelerated expansion of the universe and gravity possesses a repulsive force, At the early stage of the universe, i.e.
+Inflation corresponds to the accelerated expansion of the universe and gravity possesses a repulsive force.
+At the early stage of the universe, i.e.
 inflation, the matter is absent, $\rho=p=0$.
@@
 during inflation the Hubble parameter is a constant.
-As a result, the scale factor is given by $a(t)=a_{0}\exp(Ht)$ and corresponds to de Sitter spacetime.
-It should be noted that de Sitter stage corresponds to the eternal inflation.
-The Hubble parameter at the early universe (inflation) is $H\approx 10^{-3}M_{Pl}$ ($M_{Pl}=1/\sqrt{16\pi G}$).
-Then $GH^{2}\approx 2\cdot 10^{-8}$.
+As a result, the scale factor is given by $a(t)=a_{0}\exp(Ht)$ and corresponds to the de Sitter spacetime.
+It should be noted that de Sitter stage leads to the eternal inflation.
+The Hubble parameter at the early universe (inflation) is $H\approx 10^{-3}M_{Pl}$ ($M_{Pl}=1/\sqrt{8\pi G}$).
+Then $GH^{2}\approx 4\cdot 10^{-8}$.
 From Eq.
-(17), at $\rho=0$ we obtain the equation as follows:
-$H^{2}-\frac{b^{2}}{H^{2}+b}-2b\ln\left(\frac{H^{2}+b}{b}\right)=0.$ (32)
-Introducing new variable $x=H^{2}/b$, Eq. (32) becomes
-$x-\frac{1}{x+1}-2\ln(x+1)=0.$ (33)
-The non-trivial solution to Eq. (33) is given by $x=3.04002$ and we have the relation
+(17), at $\rho=0$ and using the variable $x=H^{2}/b$ we obtain the equation as follows:
+$x-\frac{1}{x+1}-2\ln(x+1)=0,$ (40)
+which is identical with the Eq. (29) for the phantom divide. The non-trivial solution to Eq. (40) is given by $x=3.04002$ and we have the relation
 $H^{2}=3.04002b.$ (41)
@@
 (41) and the entropy parameter $\gamma=bG/\pi$.
-Thus, we obtain $\gamma\approx 2\times 10^{-9}$.
-It is worth noting that to have a viable model of inflation one needs a smooth graceful exit into the deceleration stage [1].
+Thus, we obtain $\gamma\approx 4.2\times 10^{-9}$.
+It is worth noting that to have a viable model of inflation one needs a smooth graceful exit into the deceleration phase [1].
 5 F(T)-gravity from generalized entropy
@@
 To describe the inflationary era and the current universe accelerating expansion, one introduces a Lagrangian in the form of $F(T)$ in analogy with $F(R)$-gravity.
+The teleparallel theory of gravity based on teleparallel geometry uses the Weitzenböck connection (not the Levi-Civita connection).
+In the $F(T)$ theory the field equations are the second order which is an advantage compared with $F(R)$ theory having the fourth order equations.
+The spacetime does not have the curvature and possesses the torsion.
 The torsion scalar field $T$ is given by [45, 46]
@@
 Making use of equations (17) and (44) we find
-$\left[F(T)-2TF^{\prime}(T)\right]=-T-\frac{36b^{2}}{6b-T}-12b\ln\left(1-\frac{T^{2}}{6b}\right).$ (38)
+$F(T)-2TF^{\prime}(T)=-T-\frac{36b^{2}}{6b-T}-12b\ln\left(1-\frac{T^{2}}{6b}\right).$ (45)
 By integration of Eq. (45) we obtain the function $F(T)$:
-$F(T)=T-3\sqrt{6bT}~{}\mbox{tanh}^{-1}\sqrt{\frac{T}{6b}}$
-$-12b\ln\left(1-\frac{T}{6b}\right)+C\sqrt{T}-6b,$ (39)
-where $C$ is the integration constant. Thus, the teleparallel gravity with the function (39) corresponds to entropic cosmology with entropy (15) under consideration.
+$F(T)=T-3\sqrt{6bT}~\mbox{tanh}^{-1}\sqrt{\frac{T}{6b}}-12b\ln\left(1-\frac{T}{6b}\right)+C\sqrt{T}-6b,$ (46)
+where $C$ is the integration constant.
+It is worth noting that the term $C\sqrt{T}$ in Eq.
+(45) is the solution of the homogeneous equation $F(T)-2TF^{\prime}(T)=0$ and should be eliminated, $C=0$.
+As $T=-6H^{2}<0$, we can use the formula $i~\mbox{tanh}^{-1}(ix)=-\arctan(x)$.
+Then Eq.
+(46) is converted into the real function (at $C=0$) as follows:
+$F(T)=T+3\sqrt{-6bT}~\arctan\sqrt{-\frac{T}{6b}}-12b\ln\left(1-\frac{T}{6b}\right)-6b.$ (47)
+Some models of teleparallel gravity with different functions $F(T)$ were studied in [48, 49] Thus, the teleparallel gravity with the function (47) corresponds to entropic cosmology with entropy (15) under consideration.
 6 Conclusion
@@
 The barotropic perfect fluid and spatial flat FLRW universe were studied.
-We have obtained, from first law of apparent horizon thermodynamics, the modified Friedmann’s equations.
-The second Friedmann’s equation possesses an addition term corresponding to the density of dark energy which can be treated as a dynamical cosmological constant.
+We have obtained, from first law of apparent horizon thermodynamics, the modified Friedmann equations.
+The second Friedmann equation possesses an addition term corresponding to the density of dark energy which can be treated as a dynamical cosmological constant.
 Implying that there is no interaction between various components of cosmos, and the dark energy density and pressure $\rho_{D}$ and $p_{D}$ obeys ordinary conservation law, we computed $\rho_{D}$ and $p_{D}$.
@@
 In the model under consideration the universe can have two phases, acceleration and deceleration, that is due to holographic dark energy.
-It should be noted that Barrow and Tsallis entropies also lead to cosmology which ia due to Einstein’s equations with the dynamical cosmological constant [40].
+It should be noted that Barrow and Tsallis entropies also lead to cosmology which is due to Einstein’s equations with the dynamical cosmological constant [41].
 We computed the deceleration parameter that shows the possibility, for some model parameters, to describe the acceleration at the current era.
-The transition redshifts (when $q=0$) was calculated (given in Table I) for some parameters.
-We show (see Table II) that at $b\approx 0.1$ and $w=-2/3$ we have the deceleration parameter $q_{0}\approx-0.535$ which was observed in universe at the current era [39].
-Thus, our approach, based on new entropy and leading to modified Friedmann equations, can describe universe inflation and late time universe acceleration.
-It has been proven that entropic cosmology with our entropy proposed is equivalent to cosmology based on the teleparallel gravity with the function $F(T)$.
+We show that at the entropy parameter $\gamma\approx 0.055GH_{0}^{2}$ and w=0.283w=-0.283 we have the deceleration parameter $q_{0}\approx-0.535$ and the normalized density parameter $\Omega_{m0}\approx 0.315$ which were observed at the current era [40].
+We showed that our model possesses the phantom divide for the EoS of dark energy.
+Thus, our approach, based on new entropy and leading to modified Friedmann equations, can describe the universe inflation and the late time of universe acceleration.
+It has been proven that entropic cosmology with our entropy proposed is equivalent to cosmology based on the teleparallel gravity with the function $F(T)$ (Eq.
+(47)).